# A Symmetry-Preserving Reduced-Order Observer

### v1

@@ IV INVARIANT OBSERVER
We now aim to find sufficient conditions for the pre-observer in Theorem 1 to be a $G$-invariant reduced-order observer. That is, we seek conditions under which $\mathcal{Z}$ is asymptotically attractive. Like [2], we consider error coordinates that are $G$-invariant. Specifically, we take
@@
to be invariant coordinates that are non-zero if and only if $({z},{x},{y})\notin\mathcal{Z}$. Thus, ${\eta}\to{0}$ as $t\to\infty$ implies $\mathcal{Z}$ is asymptotically attractive. Let ${X}={\varphi}_{\gamma({y})}({x})$, ${Y}={\varrho}_{\gamma({y})}({y})$, and ${U}={\psi}_{\gamma({y})}({u})$. Using the moving frame to define these transformed points means $({X},{Y},{U})$ comprises a complete set of invariants [13, Ch. 8]. As will be shown shortly, the stability of the pre-observer (13) depends only ${\eta}$ and the invariants ${X}$, ${Y}$, and ${Z}$ (see Remark 1).
@@
${\lambda}({y};{x})={\varphi}_{\gamma({y})}({x})$
where ${x}$ is held constant. Then,
$\mathrm{T}{\lambda}({y};{x})\cdot{h}({x},{y},{u})=\mathrm{T}{\lambda}({Y};{X})\cdot{h}({X},{Y},{U})$ (16)
@@
where we again use the equivariance of the moving frame $\gamma$. Let
${\Lambda}({y};{x})={\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))$
Since ${\Lambda}$ is a composition of maps, it follows that its tangent map is
$\mathrm{T}{\Lambda}({y};{x})=\mathrm{T}{\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))\cdot\mathrm{T}{\varrho}_{g}({y})$
Then since ${\Lambda}({y};{x})={\lambda}({y};{x})$, we have
$\mathrm{T}{\lambda}({y};{x})=\mathrm{T}{\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))\cdot\mathrm{T}{\varrho}_{g}({y})$
Applying $\mathrm{T}{\lambda}({y};{x})$ to the vector field ${h}$, we obtain
$\mathrm{T}{\lambda}({y};{x})\cdot{h}({x},{y},{u})\\ =\mathrm{T}{\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))\cdot\mathrm{T}{\varrho}_{g}({y})\cdot{h}({x},{y},{u})$
Finally, by the $G$-invariance of the dynamics,
$\mathrm{T}{\lambda}({y};{x})\cdot{h}({x},{y},{u})\\ =\mathrm{T}{\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))\cdot{h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
which verifies Eq. (16). ∎
@@
$\dot{{\eta}}={f}({X}+{\eta},{Y},{U})-{f}({X},{Y},{U})\\ -\mathrm{T}{\beta}({Y})\cdot\big{(}{h}({X}+{\eta},{Y},{U})-{h}({X},{Y},{U})\big{)}\\ +\mathrm{T}{\lambda}({Y};{\eta})\cdot{h}({X},{Y},{U})$ (17)
@@
$\dot{{\eta}}=\mathrm{T}{\varphi}_{\gamma}({z}+{\beta}({y}))\cdot\big{(}{\alpha}({z},{y},{u})+\mathrm{T}{\beta}({y}))\cdot{h}({x},{y},{u})\big{)}\\ -\mathrm{T}{\varphi}_{\gamma}({x})\cdot{f}({x},{y},{u})+\mathrm{T}{\lambda}({y};{z}+{\beta}({y})-x)\cdot{h}({x},{y},{u})$
@@
$\dot{{\eta}}=\mathrm{T}{\varphi}_{\gamma}({z}+{\beta}({y}))\cdot\big{(}{f}({z}+{\beta}({y}),{y},{u})\big{)}-\mathrm{T}{\varphi}_{\gamma}({x})\cdot{f}({x},{y},{u})\\ -\mathrm{T}{\varphi}_{\gamma}({\beta}({y}))\cdot\mathrm{T}{\beta}({y})\cdot\big{(}{h}({z}+{\beta}({y}),{y},{u})-{h}({x},{y},{u})\big{)}\\ +\mathrm{T}{\lambda}({y};{z}+{\beta}({y})-x)\cdot{h}({x},{y},{u})$
@@
$\dot{{\eta}}={f}\big{(}{\varphi}_{\gamma}({z}+{\beta}({y})),{Y},{U}\big{)}-{f}\big{(}{X},{Y},{U}\big{)}\\ -\mathrm{T}{\beta}({Y})\cdot\Big{(}{h}\big{(}{\varphi}_{\gamma}({z}+{\beta}({y})),{Y},{U}\big{)}-{h}\big{(}{X},{Y},{U}\big{)}\Big{)}\\ +\mathrm{T}{\lambda}({y};{z}+{\beta}({y})-x)\cdot{h}({x},{y},{u})$
Substituting
@@
and using Lemma 2, we obtain Eq. (17). ∎
@@
The error system (17) depends only on the invariant error ${\eta}$ and the invariants ${X}$, ${Y}$, and ${U}$, which can be reduced to a set of $n+p-r$ functionally independent invariants, ${I}({x},{y},{u})$ [13, Ch. 8]. This observation is consistent with the full-order case considered in [2, Theorem 3].
@@ V EXAMPLE: RIGID-BODY VELOCITY OBSERVER
Consider a rigid aircraft instrumented with an accelerometer, gyroscope, magnetometer, and GNSS receiver such that its position, ${q}$, and attitude rotation matrix, ${R}_{\mathrm{IB}}$, are known without error. Furthermore, assume the angular velocity, ${\omega}$, and body-frame specific force, ${a}$, (obtained from filtered accelerometer readings) are available as inputs for the observer design. However, suppose that the body velocity, ${v}=(\mathrm{u,v,w})$ is not directly measured. The aim is to design a reduced-order velocity observer for the system
@@
$\mathrm{T}{\varphi}_{g}({x})\cdot{f}({x},{y},{u})$ $={R}_{g}({v}\times{\omega})+{R}_{g}{R}_{\mathrm{IB}}^{\mathsf{T}}{g}+{R}_{g}{a}$
@@
$\mathrm{T}{\varphi}_{g}({x})\cdot{h}({x},{y},{u})$ $=\begin{pmatrix}{R}_{\mathrm{IB}}{v}\\ {R}_{\mathrm{IB}}{S}({\omega}){R}_{g}^{\mathsf{T}}\end{pmatrix}$
@@
Because the transformation group is also linear in the measured part of the state, we can choose ${\ell}$ to simply be
@@
where ${L}\in\mathbb{R}^{3\times 3}$. Therefore,
@@
${\alpha}({z},{x},{y})=\underbrace{({z}+{R}_{\mathrm{IB}}^{\mathsf{T}}{L}{q})\times{\omega}+{R}_{\mathrm{IB}}^{\mathsf{T}}{g}+{a}}_{{f}({z}+{\beta}({y}),{y},{u})}\\ \underbrace{+{S}({\omega}){R}_{\mathrm{IB}}^{\mathsf{T}}{L}{q}-{R}_{\mathrm{IB}}^{\mathsf{T}}{L}{R}_{\mathrm{IB}}({z}+{R}_{\mathrm{IB}}^{\mathsf{T}}{L}{q})}_{-\mathrm{T}{\beta}({y})\cdot{h}({z}+{\beta}({y}),{y},{u})}$
@@
is asymptotically stable with respect to $\mathcal{Z}$. Therefore, if $(-{L})$ is Hurwitz, then the pre-observer $\dot{{z}}={\alpha}({z},{y},{u})$ is a globally exponentially stable, reduced-order, $\mathsf{SO}(3)$-invariant observer.
As a numerical example, consider the maneuvering flight trajectory shown in Fig. 4. For the initial condition $\hat{{v}}(0)={0}$ and gain matrix ${L}=10\mathbb{I}$, the time history of velocity estimates is shown in Fig. 5. To stress the observer, we include noisy measurements of ${y}$ and ${u}$. Specifically, suppose
@@
where ${w}_{q}$, ${w}_{R_{\mathrm{IB}}}$, ${w}_{\omega}$, and ${w}_{a}$ are zero-mean, Gaussian, continuous-time, “white noise” with power spectral densities $5\times 10^{-4}\mathbb{I}~{}\frac{\mathrm{m}^{2}}{\mathrm{Hz}}$, $10^{-7}\mathbb{I}~{}\frac{1}{\mathrm{Hz}}$, $10^{-5}\mathbb{I}~{}\frac{(\mathrm{rad}/\mathrm{s})^{2}}{\mathrm{Hz}}$, and $2\times 10^{-2}\mathbb{I}~{}\frac{(\mathrm{m}/\mathrm{s}^{2})^{2}}{\mathrm{Hz}}$, respectively. Fig. 6 shows the velocity estimates when ${y}$ and ${u}$ are corrupted by a realization of these random processes. While proof of stability for this case is beyond the scope of this paper, the results shown in Fig. 6 are indicative of the observer’s inherent robustness to disturbances, as expected from the fact that the undisturbed invariant error system is globally exponentially stable.
Figure 4: Maneuvering aircraft
Figure 5: Velocity estimates
@@ VI CONCLUSIONS
A new symmetry-preserving, reduced-order observer has been presented. This approach is beneficial when part of the system’s state is known with negligible error, avoiding unnecessary re-estimation of known signals and reducing computational complexity. Furthermore, the observer preserves symmetry. That is, the state estimate dynamics are invariant under the action of some Lie Group. As a result, the state estimate error dynamics are also invariant under the group action. Exploiting these symmetries can simplify the selection of observer parameters as seen with the example of a rigid-body velocity observer. Tuning the proposed observer consists of finding a smooth map ${\beta}$ such that the origin of the invariant error system is asymptotically stable. For some systems, this problem is reduced to choosing a gain matrix, ${L}$, as shown in the example. By leveraging the geometry of the problem, the proposed observer simplifies both the design process and stability analysis, providing a powerful tool for nonlinear systems.

### v2

@@ IV INVARIANT OBSERVER
We now aim to find sufficient conditions for the pre-observer in Theorem 1 to be a $G$-invariant reduced-order observer. That is, we seek conditions under which $\mathcal{Z}$ is asymptotically attractive. Practically, this aim is accomplished by choosing the map $\ell:\mathcal{Y}\to\mathcal{X}$ such that a stability claim can be made about the origin of a state estimation error system. Like [12], we consider nonlinear error coordinates that are $G$-invariant. Specifically, let
@@
be invariant coordinates that are non-zero if and only if $(z,x,y)\notin\mathcal{Z}$. Thus, $\eta\to{0}$ as $t\to\infty$ implies $\mathcal{Z}$ is asymptotically attractive. Let $X=\varphi_{\gamma(y)}(x)$, $Y=\varrho_{\gamma(y)}(y)$, and $U=\psi_{\gamma(y)}(u)$. Using the moving frame to define these transformed points means $(X,Y,U)$ constitutes a complete set of invariants [29, Ch. 8]. As will be shown shortly, the stability of the pre-observer (7) depends only $\eta$ and the invariants $X$, $Y$, and $Z$ (see Remark 1).
@@
$\lambda(y;\xi)=\varphi_{\gamma(y)}(\xi)$
where $\xi\in\mathcal{X}$ is held constant. Then,
$\mathrm{T}_{(y;\varphi_{g^{-1}}(\zeta))}\lambda\big{(}h(x,y,u)\big{)}\\ =\mathrm{T}_{(\varrho_{g}(y);\zeta)}\lambda\big{(}h(\varphi_{g}(x),\varrho_{g}(y),\psi_{g}(u))\big{)}$ (16)
for any $g\in G$ and $\zeta\in\mathcal{X}$.
@@
for any $x\in\mathcal{X}$, where we again use the equivariance of the moving frame $\gamma$. Thus,
$\lambda(y;\varphi_{g^{-1}}(\zeta))=\lambda(\varrho_{g}(y);\zeta)$
for any $\zeta\in\mathcal{X}$. Since $\lambda$ is a composition of maps, it follows that for any $\zeta\in\mathcal{X}$,
$\mathrm{T}_{(y;\varphi_{g^{-1}}(\zeta))}\lambda=\mathrm{T}_{(\varrho_{g}(y);\zeta)}\lambda\circ\mathrm{T}_{y}\varrho_{g}$
Applying this tangent map to the $G$-invariant vector field $h$, we obtain (16). ∎
@@
$\dot{\eta}=f(X+\eta,Y,U)-f(X,Y,U)\\ -\mathrm{T}_{Y}\beta\big{(}h(X+\eta,Y,U)-h(X,Y,U)\big{)}\\ +\mathrm{T}_{(Y;\eta)}\lambda\big{(}h(X,Y,U)\big{)}$ (17)
@@
$\dot{\eta}=\mathrm{T}_{z+\beta(y)}\varphi_{\gamma}\Big{(}\alpha(z,y,u)+\mathrm{T}_{y}\beta\big{(}h(x,y,u)\big{)}\Big{)}\\ -\mathrm{T}_{x}\varphi_{\gamma}\big{(}f(x,y,u)\big{)}+\mathrm{T}_{(y;z+\beta(y)-x)}\lambda\big{(}h(x,y,u)\big{)}$
@@
$\dot{\eta}=\mathrm{T}_{z+\beta(y)}\varphi_{\gamma}\big{(}f(z+\beta(y),y,u)\big{)}-\mathrm{T}_{x}\varphi_{\gamma}\big{(}f(x,y,u)\big{)}\\ -\big{(}\mathrm{T}_{\beta(y)}\varphi_{\gamma}\circ\mathrm{T}_{y}\beta\big{)}\big{(}h(z+\beta(y),y,u)-h(x,y,u)\big{)}\\ +\mathrm{T}_{(y;{z}+\beta(y)-x)}\lambda\big{(}h(x,y,u)\big{)}$
@@
$\dot{\eta}=f\big{(}\varphi_{\gamma}(z+\beta(y)),Y,U\big{)}-f\big{(}X,Y,U\big{)}\\ -\mathrm{T}_{Y}\beta\Big{(}h\big{(}\varphi_{\gamma}(z+\beta(y)),Y,U\big{)}-h\big{(}X,Y,U\big{)}\Big{)}\\ +\mathrm{T}_{(y;z+\beta(y)-x)}\lambda\big{(}h(x,y,u)\big{)}$
Notice the last term in the above equation can also be written as
$\mathrm{T}_{(y;\varphi_{\gamma^{-1}(y)}(\eta))}\lambda\big{(}h(x,y,u)\big{)}$
Therefore, we can use Lemma 2 along with a substitution of
@@
to obtain (17). ∎
Theorem 2 states sufficient conditions for the reduced-order pre-observer constructed in Theorem 1 to be an asymptotically stable observer. In particular, the error system (17) is $G$-invariant, meaning stability can be equivalently analyzed under arbitrary transformation by the Lie group’s action.
@@
The error system (17) depends only on the invariant error $\eta$ and the invariants $X$, $Y$, and $U$, which can be reduced to a set of n+prn+p-r functionally independent invariants, $I(x,y,u)$ [29, Ch. 8]. This observation is consistent with the full-order case considered in [12, Theorem 3].
@@ V EXAMPLE: RIGID-BODY VELOCITY OBSERVER
As an example, consider a rigid aircraft instrumented with an accelerometer, gyroscope, magnetometer, and GNSS receiver such that its position, $q$, and attitude rotation matrix, $R_{\mathrm{IB}}$, are known with negligible error. Furthermore, assume the angular velocity, $\omega$, and body-frame specific force, $a$, (obtained from filtered accelerometer readings) are available as inputs for the observer design. However, suppose that the body velocity, $v=(\mathrm{u,v,w})$ is not directly measured. The aim is to design a reduced-order velocity observer for the system
@@
$\mathrm{T}_{x}\varphi_{g}\big{(}f(x,y,u)\big{)}$ $=R_{g}(v\times\omega)+R_{g}R_{\mathrm{IB}}^{\mathsf{T}}g+R_{g}a$
@@
$\mathrm{T}_{y}\varrho_{g}\big{(}h(x,y,u)\big{)}$ $=\begin{pmatrix}R_{\mathrm{IB}}v\\ R_{\mathrm{IB}}{S}(\omega)R_{g}^{\mathsf{T}}\end{pmatrix}$
@@
Because the transformation group is also linear in the measured part of the state, we can choose $\ell$ to be
@@
where $L\in\mathbb{R}^{3\times 3}$ is a tuning parameter. Therefore,
@@
$\alpha(z,x,y)=\underbrace{(z+R_{\mathrm{IB}}^{\mathsf{T}}Lq)\times{\omega}+R_{\mathrm{IB}}^{\mathsf{T}}g+a}_{f(z+\beta(y),y,u)}\\ \underbrace{+{S}(\omega)R_{\mathrm{IB}}^{\mathsf{T}}Lq-R_{\mathrm{IB}}^{\mathsf{T}}LR_{\mathrm{IB}}(z+R_{\mathrm{IB}}^{\mathsf{T}}Lq)}_{-\mathrm{T}_{y}\beta\big{(}h(z+\beta(y),y,u)\big{)}}$
@@
is asymptotically stable. Therefore, if (L)(-L) is Hurwitz, then the pre-observer $\dot{z}=\alpha(z,y,u)$ is a globally exponentially stable, reduced-order, $\mathsf{SO}(3)$-invariant observer.
As a numerical demonstration, the aircraft equations of motion (18) were simulated using the nonlinear aerodynamic model in Simmons et al. [31] to capture the specific force $a$. Large-amplitude inputs to the aircraft’s throttle, aileron, elevator, and rudder were commanded, resulting in maneuvering flight trajectory shown in Fig. 4.
Figure 4: Maneuvering aircraft
Here, the aircraft starts in northbound, straight-and-level flight at 20 m/s with zero angle of attack; that is, $v(0)=[20~0~0]^{\mathsf{T}}$, $R_{\mathrm{IB}}(0)=\mathbb{I}$, and $\omega(0)=0$. For the initial estimate $\hat{v}(0)={0}$ and gain matrix $L=10\mathbb{I}$, the time history of velocity estimates is shown in Fig. 5.
Figure 5: Velocity estimates
To stress the observer, we include noisy measurements of $y$ and $u$. Specifically, suppose
@@
where $w_{q}$, $w_{R_{\mathrm{IB}}}$, $w_{\omega}$, and $w_{a}$ are zero-mean, Gaussian, continuous-time, “white noise” with power spectral densities $5\times 10^{-4}\mathbb{I}~\frac{\mathrm{m}^{2}}{\mathrm{Hz}}$, $10^{-7}\mathbb{I}~\frac{1}{\mathrm{Hz}}$, $10^{-5}\mathbb{I}~\frac{(\mathrm{rad}/\mathrm{s})^{2}}{\mathrm{Hz}}$, and $2\times 10^{-2}\mathbb{I}~\frac{(\mathrm{m}/\mathrm{s}^{2})^{2}}{\mathrm{Hz}}$, respectively. Figure 6 shows the velocity estimates when $y$ and $u$ are corrupted by a realization of these random processes, assuming a noise sampling rate of 1000 Hz to approximate continuous-time white noise.
@@
Note that in this simple example, a nonlinear filter (such as an extended Kalman filter) produces less noisy estimates; reduced-order observers have a tendency to amplify measurement noise. Nonetheless, the results shown in Fig. 6 are indicative of the observer’s inherent robustness to disturbances, as expected from the fact that the undisturbed invariant error system is globally exponentially stable [32, Lemma 5.1]. As explored in [33], this local robustness also ensures that the disturbance incurred by using sampled measurements of $y$ yields a bounded result. In practice, “measured” values for low-rate signals, such as position data from GNSS, should be propagated between samples using an extended Kalman predictor, for instance. While proof of stability in the presence of noise is beyond the scope of this paper, the reader is directed to [34] for an application-specific treatment illustrating how process noise might impact performance and stability analysis.
@@ VI CONCLUSIONS
A symmetry-preserving, reduced-order observer has been presented. This approach is beneficial when part of the system’s state is known with negligible error, avoiding unnecessary re-estimation of known signals and reducing computational complexity. Furthermore, the observer preserves symmetry. That is, the state estimate dynamics are invariant under the action of a Lie group. As a result, the state estimate error dynamics are also invariant under the group action. Exploiting these symmetries can simplify the selection of observer parameters as seen with the example of a rigid-body velocity observer. Tuning the proposed observer consists of finding a smooth map $\ell$ such that the origin of the invariant error system is asymptotically stable. For some systems, this problem is reduced to choosing a gain matrix, $L$, as shown in the example. By leveraging the geometry of the problem, the proposed observer simplifies both the design process and stability analysis, providing a powerful tool for safety-critical nonlinear systems.